OpenStudy (anonymous):

Determine whether the function is a polynomial function. I will post the function

OpenStudy (anonymous):

\[\sqrt[4]{x^{5}}-x^{2}\] +3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

explanation plz? fyi the +3 is part of the function

OpenStudy (anonymous):

Exponents in polynomial functions are positive integers.

OpenStudy (anonymous):

\[\frac{5}{4} \notin \mathbb{N}\]

OpenStudy (anonymous):

@Calcmathlete so it is not?

OpenStudy (anonymous):

It is not a polynomial function because polynomial functions MUST have positive integers for exponents.

OpenStudy (anonymous):

but in my book, they gave -3x^6-3x^5+18x^4 as a polynomial function and they say it is not a polynomial function if the EXPONENT on the variable is not a nonnegative integer

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (lgbasallote):

\[\sqrt[4]{x^5} = x^{5/4}\] 5/4 is NOT an integer

OpenStudy (lgbasallote):

integer are whole numbers

OpenStudy (lgbasallote):

so basically a polynomial can only have exponents like 0,1,2,3,4,.... all nonnegative whole numbers (if you want to phrase it that way) exponents

OpenStudy (lgbasallote):

btw..for the sake of information.. 5/4 is a rational number (fraction).

OpenStudy (anonymous):

ohhh thank you so much! it makes so much sense now!

OpenStudy (lgbasallote):

<tips imaginary hat>